# Cobordism of Domes Over Curves

### v1

@@ 3. Spaces of Polygons and Polyhedra
This section introduces definitions and proves technical results we need to prove Theorem 1.2, that not all integral curves are cobordant to a unit rhombus. Our approach is inspired by techniques introduced by Anan’in and Korshunov, who gave an altenate proof that Kenyon’s question is false for almost all integral curves in [1]. Specifically, we generalize Anan’in and Korshunov’s results to graph surfaces with multiple boundary components. We state our definitions here. For the original definitions and for a more detailed introduction, see [1, §2, §3].
@@
A sample polygon is a finite $1$-complex $P=(U,F)$ whose underling space is homeomorphic to $S^{1}$. Here $U$ is a set of vertices and $F=\{f_{1},\dotsc,f_{k}\}$ is a set of edges oriented and cyclically ordered with respect to the orientation of the circle. $F$ is equipped with an edge length function $\ell:F\to\mathbb{R}_{>0}$ which satisfies a nondegeneracy condition $2\ell(f_{i})<\ell(f_{1})+\dotsc+\ell(f_{n})$ for all $1\leq i\leq n$.
@@
For a sample polygon $P$, the space of polygons $\mathbb{E}^{P}$ is the set of all continuous maps $P\to\mathbb{E}$ that are isometries on edges. That is, $\mathbb{E}^{P}$ is the set of all realizations of $P$ is Euclidean space. Let $\text{Isom}^{+}\mathbb{E}$ be the group of orientation preserving isometries of $\mathbb{E}$. The moduli space of polygons $\mathbb{E}^{P}/\text{Isom}^{+}\mathbb{E}$ is the set of all realizations up to rigid motions. Let $\mathbb{E}$ be the subgroup of $\text{Isom}^{+}\mathbb{E}$ of translations, and $SO(3)$ be the subgroup of rotations. Note that $\text{Isom}^{+}\mathbb{E}\cong\mathbb{E}\rtimes SO(3)$. The scheme of polygons $\mathbb{E}^{P}/\mathbb{E}$ is the set of all realizations up to translations, but not rotations. Ultimately, we need to consider $\mathbb{E}^{P}/\text{Isom}^{+}\mathbb{E}$, but in our proof we first consider $\mathbb{E}^{P}/\mathbb{E}$ and then quotient by the action of $SO(3)$.
@@
For a vector of polygons $\mathbf{P}=(P_{1},\dotsc,P_{n})$, let $\mathbb{E}^{\textbf{P}}$ refer to the product $\mathbb{E}^{P_{1}}\times\dotsc\times\mathbb{E}^{P_{n}}$. Similarly, let $\mathbb{E}^{\text{P}}/\text{Isom}^{+}\mathbb{E}$ and $\mathbb{E}^{\textbf{P}}/\mathbb{E}$ refer to the products $\mathbb{E}^{P_{1}}/\text{Isom}^{+}\mathbb{E}\times\dotsc\times\mathbb{E}^{P_{n}}/\text{Isom}^{+}\mathbb{E}$ and $\mathbb{E}^{P_{1}}/E\times\dotsc\times\mathbb{E}^{P_{n}}/\mathbb{E}$, respectively. As an abuse of terminology, we still refer to these as the space of polygons, moduli space of polygons and scheme of polygons respectively. Note that the group of isometries always acts on each sample polygon separately, see §$\S$ 5.1. To refer to the polygons explicity, we may write $\mathbb{E}^{P_{1},\dotsc,P_{n}}$ rather than $\mathbb{E}^{\mathbf{P}}$. (See Lemma  4.2.) We refer to the sets of edges as $F_{1},\dotsc,F_{n}$, individual edges as $f^{i}_{j}\in F_{i}$, and the edge length functions as $\ell_{i}:F_{i}\to\mathbb{R}_{>0}$.
@@
As the name suggests, $\mathbb{E}^{\textbf{P}}/\mathbb{E}$ is indeed a scheme. Note, however, that it may contain singular points. For a single polygon, $p\in\mathbb{E}^{P}/\mathbb{E}$ is singular if and only if all $p(f_{i})$ are parallel. For multiple polygons, $\textbf{p}\in\mathbb{E}^{\textbf{P}}/\mathbb{E}$ is singular if and only if some $p_{i}\in\mathbb{E}^{P_{i}}/\mathbb{E}$ is singular.
@@
Let $\textbf{p}\in\mathbb{E}^{\textbf{P}}/\mathbb{E}$ be a singular point. Then $\dim T_{\textbf{p}}(\mathbb{E}^{\textbf{P}}/\mathbb{E})=m+\dim(\mathbb{E}^{\textbf{P}}/\mathbb{E})$, where $m$ is the number of $p_{i}$ which are singular in $\mathbb{E}^{P_{i}}/\mathbb{E}$.
@@
Since $\mathbb{E}^{\textbf{P}}/\mathbb{E}=\mathbb{E}^{P_{1}}/\mathbb{E}\times\dotsc\times\mathbb{E}^{P_{n}}/\mathbb{E}$, we have
$\dim T_{\textbf{p}}(\mathbb{E}^{\textbf{P}}/\mathbb{E})=\dim T_{p_{1}}(\mathbb{E}^{P_{1}}/\mathbb{E})+\dotsc+\dim T_{p_{n}}(\mathbb{E}^{P_{n}}/\mathbb{E}).$
@@
Now for a single $P_{i}$, we have $\dim T_{p_{i}}(\mathbb{E}^{P_{i}}/\mathbb{E})=1+\dim(\mathbb{E}^{P_{i}}/\mathbb{E})$ from [1, Lemma 3.2]. Combining this for all $P_{i}$ in P gives us the result. ∎
@@
Consider the following explicit description of the scheme of polygons $\mathbb{E}^{\textbf{P}}/\mathbb{E}$ for a list of polygons P. A point $\textbf{p}\in\mathbb{E}^{\textbf{P}}/\mathbb{E}$ is a collection of maps $p_{i}:F_{i}\to\mathbb{E}$ which satisfy $\langle p_{i}(f^{i}_{j}),p_{i}(f^{i}_{j})\rangle=\ell_{i}(f^{i}_{j})^{2}$ for all $i,j$ and $\sum_{j}p_{i}(f^{i}_{j})=0$ for all $i$. Similarly, for $\textbf{p}\in\mathbb{E}^{\textbf{P}}/\mathbb{E}$, a point t in the tangent space $T_{\textbf{p}}(\mathbb{E}^{\textbf{P}}/\mathbb{E})$ is a collection of maps $t_{i}:F_{i}\to\mathbb{E}$ satisfying $\langle t_{i}(f^{i}_{j}),t_{i}(f^{i}_{j})\rangle=0$ for all $i,j$ and $\sum_{j}t_{i}(f^{i}_{j})=0$ for all $i$.
@@
We generalize the definition of a graph surface given in [1] to allow for multiple boundary components as follows. A genus  $n$ graph surface $S$ is a finite $2$-dimensional simplicial complex with nondegenerate triangles and edges contained in a closed surface $\widehat{S}$ such that the complement $\mathcal{D}:=\widehat{S}\smallsetminus S$ is homeomorphic to the disjoint union of $n$ disks. This has the following data: a set of vertices $V$, a set of edges $E$ with orientation, a set of triangles $\mathcal{T}$, and a map $\Phi:E\to E$ defined as $\Phi:e\mapsto-e$ which reverses orientation.
@@
A graph surface of genus $n$ has boundary $\partial\mathcal{D}=G_{1}\cup\dotsc\cup G_{n}$, where each $G_{i}$ is a boundary component homeomorphic to $S^{1}$. Each boundary component $G_{i}$ can be decomposed into the union $g^{i}_{1}\cup\dotsc\cup g^{i}_{k_{i}}$ with a cyclic order of the edges $g^{i}_{1},\dotsc,g^{i}_{k_{i}}\in E$ such that $g^{i}_{j}$ and $g^{i}_{j+1}$ are consecutive for all $i,j$. Note that the list $g^{i}_{1},\dotsc,g^{i}_{k_{i}}$ admits repetitions with the same or opposite orientation. As an important special case, note that when a graph surface $S\subset\widehat{S}$ is contained in an orientable surface and contains no triangles, then every edge appears exactly twice, once with each orientation, in the boundary $\partial\mathcal{D}$. See [1,  Remark 2.5], and note that each boundary component can be given the same orientation with respect to the oriented surface $\widehat{S}$.
@@
Similarly, we generalize the definition of a sample polyhedron to allow for a genus $n$ graph surface instead of just a genus $1$ graph surface as in [1, Definition 2.1]. For us, a sample polyhedron will mean a genus $n$ graph surface $S$ equipped with an edge length function $\ell:\mathbb{E}\to\mathbb{R}_{>0}$ satisfying $\ell(e)=\ell(-e)$ for any edge $e\in E$ and $\ell(e_{1})+\ell(e_{2})>\ell(e_{3})$ for any triangle $T\in\mathcal{T}$ with boundary $\partial T=e_{1}+e_{2}+e_{3}$. For a given sample polyhedron $S$, we also consider the space of polyhedra $\mathbb{E}^{S}$, the moduli space of polyhedra $\mathbb{E}^{S}/\text{Isom}^{+}\mathbb{E}$ and the scheme of polyhedra $\mathbb{E}^{S}/\mathbb{E}$.
@@
An explicit description for the scheme of polyhedra $\mathbb{E}^{S}/\mathbb{E}$ is given by the set of continuous maps $q:E\to\mathbb{E}$ satisfying
@@
Similarly, an explicit description for the tangent space $T_{q}(\mathbb{E}^{S}/\mathbb{E})$ is given by the set of continuous maps $s:E\to\mathbb{E}$ satisfying
@@
    $\langle s(e),s(e)\rangle=0$ for all $e\in E$,
and the previous conditions (ii) - (iv). See [1, §§2.7-9].
@@
The boundary map induces a continuous map on schemes $\delta:\mathbb{E}^{S}/\mathbb{E}\to\mathbb{E}^{\textbf{P}}/\mathbb{E}$ and a derivative map on the tangent spaces $d\delta:T_{q}(\mathbb{E}^{S}/\mathbb{E})\to T_{q\circ\delta}(\mathbb{E}^{\textbf{P}}/\mathbb{E})$ for any point $q:E\to\mathbb{E}$ in the scheme of polyhedra $\mathbb{E}^{S}/\mathbb{E}$. These maps are from precomposition with $\delta$, see [1, §2.10].
Lastly, we define the collapse of a genus $n$ graph surface $S$. Let $T\in\mathcal{T}$ be a triangle whose boundary $\partial T$ contains a boundary edge $g^{i}_{j}$. We may collapse $S$ at $T$. The resulting simplicial complex $S^{\prime}\subset\widehat{S}$ has the same vertices $V^{\prime}=V$, one less pair of oriented edges $E^{\prime}=E\smallsetminus\{g^{i}_{j},-g^{i}_{j}\}$, and one less triangle $\mathcal{T}^{\prime}=\mathcal{T}\smallsetminus\{T\}$. Note that $S^{\prime}$ is again a genus $n$ graph surface. The boundary component $G_{i}$ changes from $g^{i}_{1}\cup\dotsc\cup g^{i}_{j}\cup\dotsc\cup g^{i}_{k_{i}}$ to $g^{i}_{1}\cup\dotsc\cup-e^{\prime}\cup-e\cup\dotsc\cup g^{i}_{k_{i}}$, where $\partial T=g^{i}_{j}+e+e^{\prime}$. All other boundary components are unchanged. In terms of a realization $q:E\to\mathbb{E}$ in the scheme of polyhedra $\mathbb{E}^{S}/\mathbb{E}$, collapse corresponds to restriction.
@@
Let $S^{\prime}$ be a sample polyhedra obtained from $S$ by collapse of a triangle, and let $q:E\to\mathbb{E}$ be a realization in the scheme of polyhedra $\mathbb{E}^{S}/\mathbb{E}$. Denote by $q^{\prime}=q|_{E^{\prime}}:E^{\prime}\to E$ the restriction of $q$ to $E^{\prime}$. Then $q^{\prime}\in\mathbb{E}^{S^{\prime}}/\mathbb{E}$. Similarly, for $s:E\to\mathbb{E}$ in $T_{q}(\mathbb{E}^{S}/\mathbb{E})$, denote by $s^{\prime}=s|_{E^{\prime}}:E^{\prime}\to\mathbb{E}$ the restriction of $s$ to $E^{\prime}$. Then $s^{\prime}\in T_{q^{\prime}}(\mathbb{E}^{S^{\prime}}/\mathbb{E})$.
@@
This is exactly the statement of [1, Proposition 4.2 (i)]. The same proof follows exactly when we allow for a genus $n$ graph surface. ∎
Consider the group $SO(3)^{\times n}$ acting on $\mathbb{E}^{\textbf{P}}/\mathbb{E}$, where each copy of $SO(3)$ acts on the corresponding scheme of polygons $\mathbb{E}^{P_{i}}/\mathbb{E}$. For $\textbf{p}\in\mathbb{E}^{\textbf{P}}/\mathbb{E}$, we need to describe the tangent space $T_{\textbf{p}}SO(3)^{\times n}\textbf{p}$. Recall the Lie algebra $\mathfrak{so}_{3}$ of the Lie group $SO(3)$ with the following description.
@@
Let $\mathfrak{so}_{3}^{\times n}$ be the set of vectors $\textbf{a}=(a_{1},\dotsc,a_{n})$, where each $a_{i}\in\mathfrak{so}_{3}$. We have the following description of the tangent space to the $SO(3)^{\times n}$ orbit on $\mathbb{E}^{\textbf{P}}/\mathbb{E}$.
@@
Let $\textbf{p}\in\mathbb{E}^{\textbf{P}}/\mathbb{E}$. The tangent space $T_{\textbf{p}}SO(3)^{\times n}\textbf{p}$ to the $SO(3)^{\times n}$-orbit of p is given by
@@
The first line is immediate because each copy of $SO(3)$ acts independently on each copy of $\mathbb{E}^{P_{i}}/\mathbb{E}$. Then for a single $SO(3)$ acting on $p_{i}\in\mathbb{E}^{P_{i}}/\mathbb{E}$, we know $T_{p_{i}}SO(3)p_{i}=\{a_{i}\circ p_{i}:F_{i}\to\mathbb{E}\ |\ a_{i}\in\mathfrak{so}_{3}\}$ from [1, §3.1]. ∎
The final ingredient is a symplectic form on the scheme of polygons $\mathbb{E}^{\textbf{P}}/\mathbb{E}$. Consider the skew symmetric form $\omega$ given by the formula
@@
The tangent space $T_{\textbf{p}}(SO(3)^{\times n}\textbf{p})$ to the $SO(3)^{\times n}$ orbit of any point $\textbf{p}\in\mathbb{E}^{\textbf{P}}/\mathbb{E}$ coincides with the kernel of the form $\omega_{\textbf{p}}$ on $T_{\textbf{p}}(\mathbb{E}^{\textbf{P}}/\mathbb{E})$.
@@
For a single $p_{i}\in\mathbb{E}^{P_{i}}/\mathbb{E}$, we know the kernel of the form $\omega_{p_{i}}$ on $T_{p_{i}}(\mathbb{E}^{P_{i}}/\mathbb{E})$ corresponds to $T_{p_{i}}SO(3)p_{i}$ from [1, Lemma  3.4]. This immediately shows that $T_{\textbf{p}}SO(3)^{\times n}\textbf{p}\subset\ker\omega_{\textbf{p}}$, as each term in the sum will be 00. For the reverse inclusion, we repeat the dimension counting argument from [1,  Lemma 3.4] with our Lemma 3.1 in place of Lemma [1,  Lemma 3.2]. This shows the spaces have the same dimension, so they are equal. ∎

### v2

@@ 3. Spaces of Polygons and Polyhedra
This section introduces definitions and proves technical results we need to prove Theorem 4.3, that not all integral curves are orientably cobordant to a unit rhombus. Our approach is inspired by techniques introduced by Anan’in and Korshunov, who gave an alternate proof of the negative answer to Kenyon’s question for almost all integral curves in [AK24]. Specifically, we generalize Anan’in and Korshunov’s results to graph surfaces with multiple boundary components corresponding to the union of multiple integral curves. We state our definitions here. For the original definitions and for a more detailed introduction, see [AK24, §2, §3].
@@
A sample polygon is a finite $1$-complex $P=(U,F)$ whose underlying space is homeomorphic to $S^{1}$. Here $U$ is a set of vertices and $F=\{f_{1},\dotsc,f_{k}\}$ is a set of edges oriented and cyclically ordered with respect to the orientation of the circle. $F$ is equipped with an edge length function $\ell:F\to\mathbb{R}_{>0}$ which satisfies a nondegeneracy condition $2\ell(f_{i})<\ell(f_{1})+\dotsc+\ell(f_{n})$ for all $1\leq i\leq n$.
@@
For a sample polygon $P$, the space of polygons $\mathbb{E}^{P}$ is the set of all continuous maps $p:P\to\mathbb{E}$ that are isometries on edges. That is, $\mathbb{E}^{P}$ is the set of all realizations of $P$ in Euclidean space. Let $\text{Isom}^{+}\mathbb{E}$ be the group of orientation preserving isometries of $\mathbb{E}$. The moduli space of polygons $\mathbb{E}^{P}/\text{Isom}^{+}\mathbb{E}$ is the set of all realizations up to rigid motions. Let $\mathcal{E}$ be the subgroup of $\text{Isom}^{+}\mathbb{E}$ of translations, and $SO(3)$ be the subgroup of rotations. Note that $\text{Isom}^{+}\mathbb{E}\cong\mathcal{E}\rtimes SO(3)$. The scheme of polygons $\mathbb{E}^{P}/\mathcal{E}$ is the set of all realizations up to translations, but not rotations. Ultimately, we need to consider $\mathbb{E}^{P}/\text{Isom}^{+}\mathbb{E}$, but in our proof we first consider $\mathbb{E}^{P}/\mathcal{E}$ and then quotient by the action of $SO(3)$.
@@
For a vector of polygons $\mathbf{P}=(P_{1},\dotsc,P_{n})$, let $\mathbb{E}^{\textbf{P}}$ refer to the product $\mathbb{E}^{P_{1}}\times\dotsc\times\mathbb{E}^{P_{n}}$. Similarly, let $\mathbb{E}^{\textbf{P}}/\text{Isom}^{+}\mathbb{E}$ and $\mathbb{E}^{\textbf{P}}/\mathcal{E}$ refer to the products $\mathbb{E}^{P_{1}}/\text{Isom}^{+}\mathbb{E}\times\dotsc\times\mathbb{E}^{P_{n}}/\text{Isom}^{+}\mathbb{E}$ and $\mathbb{E}^{P_{1}}/\mathcal{E}\times\dotsc\times\mathbb{E}^{P_{n}}/\mathcal{E}$, respectively. As an abuse of terminology, we still refer to these as the space of polygons, moduli space of polygons and scheme of polygons respectively. Note that the group of isometries always acts on each sample polygon separately, see §$\S$ 5.1. To refer to the polygons explicity, we may write $\mathbb{E}^{P_{1},\dotsc,P_{n}}$ rather than $\mathbb{E}^{\mathbf{P}}$. (See Lemma  4.2.) We refer to the sets of edges as $F_{1},\dotsc,F_{n}$, individual edges as $f^{i}_{j}\in F_{i}$, and the edge length functions as $\ell_{i}:F_{i}\to\mathbb{R}_{>0}$.
@@
As the name suggests, $\mathbb{E}^{\textbf{P}}/\mathcal{E}$ is indeed a scheme. Note, however, that it may contain singular points. For one polygon, $p\in\mathbb{E}^{P}/\mathcal{E}$ is singular if and only if all $p(f_{i})$ are parallel. For multiple polygons, $\textbf{p}\in\mathbb{E}^{\textbf{P}}/\mathcal{E}$ is singular if and only if some $p_{i}\in\mathbb{E}^{P_{i}}/\mathcal{E}$ is singular.
@@
Let $\textbf{p}\in\mathbb{E}^{\textbf{P}}/\mathcal{E}$ be a singular point. Then $\dim T_{\textbf{p}}(\mathbb{E}^{\textbf{P}}/\mathcal{E})=m+\dim(\mathbb{E}^{\textbf{P}}/\mathcal{E})$, where $m$ is the number of $p_{i}$ which are singular in $\mathbb{E}^{P_{i}}/\mathcal{E}$.
@@
Since $\mathbb{E}^{\textbf{P}}/\mathcal{E}=\mathbb{E}^{P_{1}}/\mathcal{E}\times\dotsc\times\mathbb{E}^{P_{n}}/\mathcal{E}$, we have
$\dim T_{\textbf{p}}(\mathbb{E}^{\textbf{P}}/\mathcal{E})=\dim T_{p_{1}}(\mathbb{E}^{P_{1}}/\mathcal{E})+\dotsc+\dim T_{p_{n}}(\mathbb{E}^{P_{n}}/\mathcal{E}).$
@@
Now for a singular $p_{i}$, we have $\dim T_{p_{i}}(\mathbb{E}^{P_{i}}/\mathcal{E})=1+\dim(\mathbb{E}^{P_{i}}/\mathcal{E})$ from [AK24, Lemma 3.2], else for a smooth $p_{i}$, we have $\dim T_{p_{i}}(\mathbb{E}^{P_{i}}/\mathcal{E})=\dim(\mathbb{E}^{P_{i}}/\mathcal{E})$. Combining this for all $p_{i}$ in P gives us the result. ∎
@@
Consider the following explicit description of the scheme of polygons $\mathbb{E}^{\textbf{P}}/\mathcal{E}$ for a list of polygons P. A point $\textbf{p}\in\mathbb{E}^{\textbf{P}}/\mathcal{E}$ is a collection of maps $p_{i}:F_{i}\to\mathbb{E}$ which satisfy $\langle p_{i}(f^{i}_{j}),p_{i}(f^{i}_{j})\rangle=\ell_{i}(f^{i}_{j})^{2}$ for all $i,j$ and $\sum_{j}p_{i}(f^{i}_{j})=0$ for all $i$. Similarly, for $\textbf{p}\in\mathbb{E}^{\textbf{P}}/\mathcal{E}$, a point t in the tangent space $T_{\textbf{p}}(\mathbb{E}^{\textbf{P}}/\mathcal{E})$ is a collection of maps $t_{i}:F_{i}\to\mathbb{E}$ satisfying $\langle t_{i}(f^{i}_{j}),p_{i}(f^{i}_{j})\rangle=0$ for all $i,j$ and $\sum_{j}t_{i}(f^{i}_{j})=0$ for all $i$.
@@
We generalize the definition of a graph surface given in [AK24] to allow for multiple boundary components as follows. A genus  $n$ graph surface $S$ is a finite $2$-dimensional simplicial complex with nondegenerate triangles and edges contained in a closed surface $\widehat{S}$ such that the complement $\mathcal{D}:=\widehat{S}\smallsetminus S$ is homeomorphic to the disjoint union of $n$ disks. This has the following data: a set of vertices $V$, a set of edges $E$ with orientation, a set of triangles $\mathcal{T}$, and a map $\Phi:E\to E$ defined as $\Phi:e\mapsto-e$ which reverses orientation.
@@
A graph surface of genus $n$ has boundary $\partial\mathcal{D}=G_{1}\cup\dotsc\cup G_{n}$, where each $G_{i}$ is a boundary component homeomorphic to $S^{1}$. Each boundary component $G_{i}$ can be decomposed into the union $g^{i}_{1}\cup\dotsc\cup g^{i}_{k_{i}}$ with a cyclic order of the edges $g^{i}_{1},\dotsc,g^{i}_{k_{i}}\in E$ such that $g^{i}_{j}$ and $g^{i}_{j+1}$ are consecutive for all $i,j$. Note that the list $g^{i}_{1},\dotsc,g^{i}_{k_{i}}$ admits repetitions with the same or opposite orientation. As an important special case, note that when a graph surface $S\subset\widehat{S}$ is contained in an orientable surface and contains no triangles, then every edge appears exactly twice, once with each orientation, in the boundary $\partial\mathcal{D}$. See [AK24,  Remark 2.5], and note that each boundary component can be given the same orientation with respect to the oriented surface $\widehat{S}$.
@@
Similarly, we generalize the definition of a sample polyhedron to allow for a genus $n$ graph surface instead of just a genus $1$ graph surface as in [AK24, Definition 2.1]. For us, a sample polyhedron will mean a genus $n$ graph surface $S$ equipped with an edge length function $\ell:\mathbb{E}\to\mathbb{R}_{>0}$ satisfying $\ell(e)=\ell(-e)$ for any edge $e\in E$ and $\ell(e_{1})+\ell(e_{2})>\ell(e_{3})$ for any triangle $T\in\mathcal{T}$ with boundary $\partial T=e_{1}+e_{2}+e_{3}$. For a given sample polyhedron $S$, we also consider the space of polyhedra $\mathbb{E}^{S}$, the moduli space of polyhedra $\mathbb{E}^{S}/\text{Isom}^{+}\mathbb{E}$ and the scheme of polyhedra $\mathbb{E}^{S}/\mathcal{E}$.
@@
An explicit description for the scheme of polyhedra $\mathbb{E}^{S}/\mathcal{E}$ is given by the set of continuous maps $q:E\to\mathbb{E}$ satisfying
@@
Similarly, an explicit description for the tangent space $T_{q}(\mathbb{E}^{S}/\mathcal{E})$ is given by the set of continuous maps $s:E\to\mathbb{E}$ satisfying
@@
    $\langle s(e),q(e)\rangle=0$ for all $e\in E$,
and the previous conditions (ii) - (iv). See [AK24, §§2.7-9].
@@
The boundary map induces a continuous map on schemes $\delta:\mathbb{E}^{S}/\mathcal{E}\to\mathbb{E}^{\textbf{P}}/\mathcal{E}$ and a derivative map on the tangent spaces $d\delta:T_{q}(\mathbb{E}^{S}/\mathcal{E})\to T_{q\circ\delta}(\mathbb{E}^{\textbf{P}}/\mathcal{E})$ for any point $q:E\to\mathbb{E}$ in the scheme of polyhedra $\mathbb{E}^{S}/\mathcal{E}$. These maps are from precomposition with $\delta$, see [AK24, §2.10].
Lastly, we define the collapse of a genus $n$ graph surface $S$. Let $T\in\mathcal{T}$ be a triangle whose boundary $\partial T$ contains a boundary edge $g^{i}_{j}$. We may collapse $S$ at $T$. The resulting simplicial complex $S^{\prime}\subset\widehat{S}$ has the same vertices $V^{\prime}=V$, one less pair of oriented edges $E^{\prime}=E\smallsetminus\{g^{i}_{j},-g^{i}_{j}\}$, and one less triangle $\mathcal{T}^{\prime}=\mathcal{T}\smallsetminus\{T\}$. Note that $S^{\prime}$ is again a genus $n$ graph surface. The boundary component $G_{i}$ changes from $g^{i}_{1}\cup\dotsc\cup g^{i}_{j}\cup\dotsc\cup g^{i}_{k_{i}}$ to $g^{i}_{1}\cup\dotsc\cup-e^{\prime}\cup-e\cup\dotsc\cup g^{i}_{k_{i}}$, where $\partial T=g^{i}_{j}+e+e^{\prime}$. All other boundary components are unchanged. In terms of a realization $q:E\to\mathbb{E}$ in the scheme of polyhedra $\mathbb{E}^{S}/\mathcal{E}$, collapse corresponds to restriction.
@@
Let $S^{\prime}$ be a sample polyhedra obtained from $S$ by collapse of a triangle, and let $q:E\to\mathbb{E}$ be a realization in the scheme of polyhedra $\mathbb{E}^{S}/\mathcal{E}$. Denote by $q^{\prime}=q|_{E^{\prime}}:E^{\prime}\to E$ the restriction of $q$ to $E^{\prime}$. Then $q^{\prime}\in\mathbb{E}^{S^{\prime}}/\mathcal{E}$. Similarly, for $s:E\to\mathbb{E}$ in $T_{q}(\mathbb{E}^{S}/\mathcal{E})$, denote by $s^{\prime}=s|_{E^{\prime}}:E^{\prime}\to\mathbb{E}$ the restriction of $s$ to $E^{\prime}$. Then $s^{\prime}\in T_{q^{\prime}}(\mathbb{E}^{S^{\prime}}/\mathcal{E})$.
@@
This is exactly the statement of [AK24, Proposition 4.2 (i)]. The same proof follows exactly when we allow for a genus $n$ graph surface. ∎
Consider the group $SO(3)^{\times n}$ acting on $\mathbb{E}^{\textbf{P}}/\mathcal{E}$, where each copy of $SO(3)$ acts on the corresponding scheme of polygons $\mathbb{E}^{P_{i}}/\mathcal{E}$. For $\textbf{p}\in\mathbb{E}^{\textbf{P}}/\mathcal{E}$, we need to describe the tangent space $T_{\textbf{p}}SO(3)^{\times n}\textbf{p}$. Recall the Lie algebra $\mathfrak{so}_{3}$ of the Lie group $SO(3)$ with the following description.
@@
Let $\mathfrak{so}_{3}^{\times n}$ be the set of vectors $\textbf{a}=(a_{1},\dotsc,a_{n})$, where each $a_{i}\in\mathfrak{so}_{3}$. We have the following description of the tangent space to the $SO(3)^{\times n}$ orbit on $\mathbb{E}^{\textbf{P}}/\mathcal{E}$.
@@
Let $\textbf{p}\in\mathbb{E}^{\textbf{P}}/\mathcal{E}$. The tangent space $T_{\textbf{p}}SO(3)^{\times n}\textbf{p}$ to the $SO(3)^{\times n}$-orbit of p is given by
@@
The first line is immediate because each copy of $SO(3)$ acts independently on each copy of $\mathbb{E}^{P_{i}}/\mathcal{E}$. Then for a single $SO(3)$ acting on $p_{i}\in\mathbb{E}^{P_{i}}/\mathcal{E}$, we know $T_{p_{i}}SO(3)p_{i}=\{a_{i}\circ p_{i}:F_{i}\to\mathbb{E}\ |\ a_{i}\in\mathfrak{so}_{3}\}$ from [AK24, §3.1]. ∎
The final ingredient is a symplectic form on the scheme of polygons $\mathbb{E}^{\textbf{P}}/\mathcal{E}$. Consider the skew symmetric form $\omega$ given by the formula
@@
The tangent space $T_{\textbf{p}}(SO(3)^{\times n}\textbf{p})$ to the $SO(3)^{\times n}$ orbit of any point $\textbf{p}\in\mathbb{E}^{\textbf{P}}/\mathcal{E}$ coincides with the kernel of the form $\omega_{\textbf{p}}$ on $T_{\textbf{p}}(\mathbb{E}^{\textbf{P}}/\mathcal{E})$.
@@
For a single $p_{i}\in\mathbb{E}^{P_{i}}/\mathcal{E}$, we know the kernel of the form $\omega_{p_{i}}$ on $T_{p_{i}}(\mathbb{E}^{P_{i}}/\mathcal{E})$ corresponds to $T_{p_{i}}SO(3)p_{i}$ from [AK24, Lemma  3.4]. This immediately shows that $T_{\textbf{p}}SO(3)^{\times n}\textbf{p}\subset\ker\omega_{\textbf{p}}$, as each term in the sum will be 00. For the reverse inclusion, we repeat the dimension counting argument from [AK24,  Lemma 3.4] with our Lemma 3.1 in place of Lemma [AK24,  Lemma 3.2]. This shows the spaces have the same dimension, so they are equal. ∎